OpenStudy (anonymous):

differentiate with respect to x 4e^cos2x

geerky42 (geerky42):

Chain rule.

geerky42 (geerky42):

\[\Large \dfrac{d}{dx}e^u = u'e^u\]

geerky42 (geerky42):

Do you know chain rule, @Nutty ?

OpenStudy (anonymous):

Ya

geerky42 (geerky42):

Well, use it. Can you do it?

OpenStudy (anonymous):

You'll have to use the chain rule twice on this one.

OpenStudy (anonymous):

Yaa

OpenStudy (anonymous):

let u=cos 2x, then\[4e^{\cos 2x}=4e^u\]

OpenStudy (anonymous):

so,\[\frac{d}{dx}4e^u=4u'e^u\]

OpenStudy (anonymous):

The only thing you need to figure out is u' What is u'?